OpenStudy (jordanloveangel):

Solve the equation for y. 8x – 9y = 11

OpenStudy (anonymous):

8x – 9y = 11 -9y=11-8x divide by -9 y=-11/9 +8/11 x

OpenStudy (greencat):

^^ Ooh It'salready solved. :D

OpenStudy (greencat):

Whoops u made a mistake.

OpenStudy (greencat):

8x-9y=11 8x=9y+11 8x-11=9y 8/9x-11/9=y so \[y=\frac{ 8 }{ 9 }x-\frac{ 11 }{ 9 }\]

OpenStudy (jordanloveangel):

oh okay thnx green <3

OpenStudy (anonymous):

if you pick a single variable which you will eliminate in further steps say X, so now transfer the X term to LHS (left hand side) that will be => 9y=8X-11

OpenStudy (greencat):

Ya lol she prbly gets it by now. she was prbly lazy.

OpenStudy (greencat):

Smiles, xtrmey@ :D